The probability density function of a random variable X is given by :

f(x)={kx(2-x), 0<x<2

{0, e.w

a) Find the value of k

b)Find the distribution function F(X)

c)Find P(X>3)


1
Expert's answer
2021-01-22T03:26:42-0500

(a) Probability density function of a random veriable "X" is given by

"f(x)=kx(2-x)" , "0<x<2"

"=0," e.w

We know that for "p.d.f" ,

"\int _{-\infty}^{\infty} f(x)dx=1"

"\implies \int _ {0}^{2} kx(2-x)dx=1"

"\implies k.\int _ {0}^{2}(2x-x^2)dx=1"

"\implies k.[2.\frac{x^2}{2}-\frac {x^3}{3}]_{0}^{2}=1"

"\implies k.[4-\frac{8}{3}]=1"

"\implies k=\frac{3}{4}"


(b) The distribution function of the continuous random variable "X" having probability density function "f(x)" is given by ,

"F_X(x)=P(X\leq x)=\int _{-\infty}^{x}f(u) du"

Therefore for the given "p.d.f" , the distribution function "F_X(x)=\intop_{-\infty}^{x}f(u)du"

"=\intop_{0}^{x}\frac{3}{4}u(2-u)du"

"=\frac{3}{4}\intop_{0}^{x}(2u-u^2)du"

"=\frac{3}{4}[u^2-\frac{u^3}{3}]_{0}^{x}"

"=\frac{1}{4}[3x^2-x^3]"


(c) "P(X>3)=\intop_{3}^{\infty}f(x)dx=0" [ as "f(x)=0" when "3<x<\infty" ]